What is [H3O^+] of the strongest acid?
a. 1E0 moles/liter
b. 1E-1 moles/liter
c. 1E1 moles/liter
d. 1E-2 moles/liter

Answer by jsmallt9(3759)

a. 1E0 moles/liter
1E0 = 1*10^0 = 1*1 = 1
b. 1E-1 moles/liter
1E-1 = 1*10^(-1) = 1*(1/10) = 1/10
c. 1E1 moles/liter
1E1 = 1*10^1 = 1*10 = 10
d. 1E-2 moles/liter
1E-2 = 1*10^-2 = 1*(1/100) = 1/100
(c) wins!